OpenStudy (anonymous):

A function has the rule "multiply by two, and then subtract one." What is the output when the input is 7? 4 twelve thirteen fifteen

OpenStudy (curry):

14

OpenStudy (anonymous):

14 is the wrong answer. The answer is 13. Why? b/c when you translate this into an equation you get: f(x)=2x-1. Now plug in 7, and you get: f(7)=2(7)-1=14-1=13